OpenStudy (anonymous):

Solve the following system of linear equations by the elimination method. 13x+3y=32 3x-2y=2

OpenStudy (zepp):

Can you multiply the first one by 2 and the second one by 3?

OpenStudy (anonymous):

let me try, hang on

OpenStudy (anonymous):

no that doesnt work:[

OpenStudy (zepp):

Alright there: (13x+3y=32)*2 (13x+3y)*2 = 32*2 13x*2 + 3y*2 = 64 26x + 6y = 64 Now try for the second one with 3 ;D

OpenStudy (anonymous):

the second one would be 9x-18y=6 now what do i do?

OpenStudy (zepp):

Nope, it would be (3x-2y)*3=2*3 9x - 6y = 6

OpenStudy (anonymous):

oh your right, wow im dumb today haha. thank you! I just have one more problem i need help with:

OpenStudy (anonymous):

-2/5x-1/2y=-1/3 -8/5x-2y=-4/3

OpenStudy (zepp):

26x + 6y = 64 9x - 6y = 6 You have this 6y cancels out 35x = 70 x = 2

OpenStudy (anonymous):

yeah i got the answer, thank you!

OpenStudy (zepp):

Okay, you have -2/5x - 1/2y = -1/3 It's equal to 2/5x + 1/2y = 1/3 Now you have 2/5x + 1/2y = 1/3 -8/5x - 2y = -4/3 +1/2 and -2y could be cancelled out, but you have to do something so in the first equation, 1/2y turns into a 2y

OpenStudy (anonymous):

so y would equal 2 and so what would x equal?

OpenStudy (zepp):

You plug the y(2) into any equation and solve for x :)

OpenStudy (anonymous):

ok yeah, i just did that and i got up to this point: 4/5x+2y=1/3, now what? im so bad with fractions

OpenStudy (zepp):

2/5x + 1/2y = 1/3 1/2 = 0.5 2/0.5 = 4 So you have multiply everything by 4 (2/5x + 1/2y)*4 = (1/3)*4 8/5x + 2y = 4/3 You have now 8/5x + 2y = 4/3 -8/5x - 2y= -4/3 Solve :)

OpenStudy (anonymous):

so x is no solution?

OpenStudy (zepp):

Umm, lemme see

OpenStudy (zepp):

x would be 1

OpenStudy (anonymous):

and y would be 2?

OpenStudy (zepp):

It would be 2/3

OpenStudy (zepp):

owait, I don't get it this one o.o

OpenStudy (anonymous):

haha its okay, thanks anyways!

OpenStudy (zepp):

@apoorvk